# Safe Learning for Contact-Rich Robot Tasks: A Survey from Classical Learning-Based Methods to Safe Foundation Models

### v1

@@ 2.2.2 Stability and Invariance: Lyapunov and CBF Views
A second axis treats safety as stability and set invariance. Lyapunov methods certify convergence or boundedness and can imply safety when safe sets are invariant sublevel sets of a Lyapunov function [khalil2002nonlinear, kumar_safe_2025]. Control Barrier Functions (CBFs) enforce forward invariance of safety sets online, minimally modifying actions to prevent constraint violations; they have seen increasing use in contact-aware robotics and safe RL [ames2019cbf, guerrier2024learning, 213, 235]. End-to-end safe RL can embed barrier-based constraints into training or runtime filtering [287, 129]. Here, safety means remaining inside certified invariant sets while preserving task feasibility.
@@ 2.2.3 Constraint Satisfaction and Reachability
Safety can be defined as the satisfaction of state, input, and contact constraints (e.g., joint/velocity/torque limits, friction cones, force envelopes). Safe MPC enforces constraints over a horizon with robust tubes or uncertainty sets [hewing2020learning, 21]. Hamilton–Jacobi reachability and safety filters compute (or approximate) safe sets and veto actions that would exit them [bansal2017hamilton, wabersich2023data, 271]. In the constraint-centric view, safety is the persistent satisfaction of physically meaningful bounds throughout approach, impact, sliding, and re-grasp phases.
@@ 2.2.4 Risk- and Data-Driven Safety
Data-driven definitions quantify safety via risk-sensitive objectives, recovery zones, or auxiliary critics that predict and penalize unsafe outcomes. Recovery RL separates exploration from recovery to reduce unsafe excursions [91]; safety critics or pessimistic objectives downweight risky actions under uncertainty [185, gu2024review, brunke2022safe]. Recent work provides theoretical formulations (e.g., safe value functions, shielding and generalization guarantees) to connect statistical learning with formal safety [165, 137]. In this axis, safety is a probabilistic guarantee or risk bound over contact outcomes under imperfect models and noisy sensing.
@@ 2.2.5 Human-Centered and Standard-Driven Safety
In human–robot collaboration and assistance, safety must address physical limits (contact pressure, speed-and-separation, ergonomics) and perceived safety (predictability, comfort, trust) [ajoudani2018progress, 52]. Standards such as ISO 10218 and ISO/TS 15066 codify limits and processes, shaping acceptable forces, speeds, and safety-rated monitoring [280, lakshminarayanan2024robots]. Perception- and policy-level safeguards further bound human exposure to risk in collaborative and assistive tasks [zhao2021efficientHRC, el2020towardsICRA, 57, 9]. Along this axis, safety extends beyond robot/object protection to human well-being and acceptability.
2.2.6 Semantic Safety with Vision–Language(–Action) Models
@@
Semantic safety concerns aligning task intent, environment context, and contact constraints with what the robot actually executes, so that language-/vision-conditioned policies do not trigger unsafe contact behaviors. In contact-rich settings, this means grounding task semantics (objects, affordances, no-touch zones, insertion axes, force limits) and safety rules (limits, PPE, human proximity) into actionable constraints and monitors for execution [yuan2022safe].
@@
Modern VLM/VLA systems can (i) parse natural-language goals and safety rules, (ii) ground them in the sensed scene (objects, surfaces, grasp points) [10705419, 33, 51], and (iii) parameterize low-level compliant controllers or skill libraries with safety-aware references (pose/force/impedance) [330, zhou_physvlm_2025, wei_audio-vla_2025] rather than direct torques. Typical architectures use a hierarchical split: a language–vision planner selects certified skills/options; mid-level skills expose safety-relevant parameters; low-level safety enforcers (impedance, CBF/MPC, rate limiters) guarantee contact stability and constraint satisfaction during execution [wei_audio-vla_2025, cui_end--end_nodate, lu_vla-rl_2025, 376].
2.3 Problem Formulation and Learning Paradigm Overview
2.3.1 Problem Formulation
@@
Contact-rich manipulation arises across assembly, insertion, forceful surface interaction, and human–robot collaboration. We consider a general formulation that does not commit to a single mathematical model. Tasks are described by (i) a state representation $x_{t}$ that includes robot kinematics/dynamics, latent contact modes, and measured or estimated wrenches; (ii) a control signal $u_{t}$ that can be low-level torques/velocities or higher-level references such as force targets and impedance/admittance gains; and (iii) safety and feasibility specifications involving force/torque envelopes, friction and approach cones, motion limits, and human-centric constraints [ajoudani2018progress, calanca2015review, yuan2022safe]. Depending on assumptions and requirements, the underlying decision model may be an unconstrained Markov decision process (MDP) or a partially observable Markov decision process (POMDP), a constrained variant (e.g., constrained Markov decision process (CMDP)), or a hybrid-systems/control-theoretic problem with explicit contact transitions. In all cases, safety is operationalized through limits on contact forces/pressures and motion, potentially augmented by risk- or semantics-derived rules [yuan2022safe, 33, 100, 330].
@@
Observability and abstraction are key design choices. When direct force/torque sensing is available, controllers can track force references and tighten safety envelopes; otherwise, observers or learned estimators provide wrench/contact-state cues for policy inputs and online monitoring [calanca2015review]. Action abstraction often improves safety and sim-to-real transfer: outputting pose/force targets or impedance parameters instead of raw torques yields compliant interaction and clearer hooks for runtime safety mechanisms [calanca2015review, ajoudani2018progress, yuan2022safe]. Finally, semantics (VLM/VLA) supply high-level task structure and safety rules that must be grounded into formal constraints and checked during execution [33, 100, 330, geng2025roboverse, gu2025robust].
@@ 2.3.2 Learning Paradigm
We summarize four main learning paradigms and how they interface with safety in contact-rich manipulation.
@@
RL: RL learns policies for variable-impedance forceful skills and insertions, often using action spaces that output force targets or stiffness/damping to preserve compliant interaction. Safety is enhanced by conservative objectives, recovery strategies, and shielded execution (e.g., barrier/reachability/MPC layers) [zhang2024srl, roveda2020model, stan2020reinforcement, zhang2025passivity, 348, 320]
@@
IL: Demonstrations initialize contact sequencing and force profiles (approach, touch, slide) and are commonly refined by RL or supervised updates; hybrid trajectory–force learning and one/few-shot strategies are prevalent in insertion and assembly settings. IL shares the same safety-friendly action abstractions and leverages compliant controllers during execution [246, 27, luu2025manifeel]
@@
Learning for Model-based and control-theoretic methods Impedance/admittance regulation, robust/adaptive force controllers, MPC with contact constraints, and barrier/reachability filters operationalize limits on contact force/torque, friction cones, and approach geometry. Learning augments these by identifying dynamics, adapting gains, or shaping feasible sets, yielding transparent enforcement at runtime [242, 235, 318, pang2023planning, 242, xu_look_2021]
@@
Foundation model-based methods Foundation models for robotics leverage large-scale VLA pretraining to parse open-ended goals, infer affordances, and impose safety-aware constraints before execution. In contact-rich settings, these methods are typically embedded in a plan– parameterize– enforce stack: (i) a VLM/VLA module interprets natural-language goals and scene context to propose task sketches and safety rules (e.g., force caps, no-touch regions, speed-and-separation); (ii) the semantic plan is compiled into formal constraints and parameterized skill references (pose/force/impedance) rather than raw torques to preserve compliant interaction; and (iii) low-level safety layers (e.g., MPC/CBF filters, rate limiters) execute with verifiable bounds on contact forces and approach geometry [33, 100, 330]. Data and evaluation rely on open toolchains and safety-oriented benchmarks to quantify violation rates, refusal/recovery behavior, and generalization under distribution shift [towers2024gymnasium, 152, geng2025roboverse, gu2025robust]. Practically, safety guardrails for foundation-model pipelines include structured prompting and tool schemas, text-to-constraint grounding with pre-execution checks, conservative default limits (force/velocity caps), and “plan-then-filter” action projection through certified controllers; these mitigations address hallucination and mis-grounding that could otherwise lead to overloads at contact [33, 330, 10705419].
@@
Hybrid integration (plan–learn–enforce) Planning and skill graphs (potentially guided by VLM/VLA) provide structure; learned policies output pose/force/impedance parameters; safety layers project actions to satisfy certified constraints. Semantic modules further align high-level intent and safety rules with the certified execution stack [geng2025roboverse, gu2025robust, 330, 10705419]. Evaluation commonly reports violations, near misses, and generalization using standardized suites and simulators [towers2024gymnasium, raffin2021stable, 152].
3 Methods Categorization from Different Perspectives
@@ 3.1 Safe Learning for Contact-Rich Tasks
3.1.1 Safe Exploration (safe learning before contact)
@@
Safe HRC has traditionally been approached through path planning algorithms such as rapidly-exploring random tree star (RRT) [chi2017risk] for dynamic human-robot systems, or artificial potential field (APF) [rodriguez2023advanced] and virtual fixtures [bowyer2013active] for real-time adjustment of the robot’s trajectory. However, these methods often yield slow trajectory computation and limited generalization, since they require task-specific tuning, assume simplified environments, and adapt poorly to dynamic or unstructured settings [bowyer2013active].
Recent approaches enforce safety constraints via control barrier functions (CBFs), where quadratic programs compute minimal trajectory deviations to maintain safety [hsu2023safety, ferraguti2020safety, dai2023safe]. Extensions to uncertain CBFs (U-CBFs) handle environmental uncertainty [das2024robust, li2023robust], though worst-case assumptions often yield conservative designs. Stochastic frameworks using chance constraints offer less conservative alternatives [wang2024flow, du2011probabilistic], but remain computationally intractable for most HRC applications due to non-convex feasible regions [ahmed2008solving]. Complementary work examines risk factors in HRC [berx2022identification, inam2018risk, liu2020dynamic, majumdar2020should]. While these approaches can better inform the tolerable risk-bound for different applications, a risk tunable CBF can solve the limitation [11023035].
@@
Reachability-based Trajectory Safeguard (RTS) [Shao2021RTS] is an offline-online safety layer that precomputes reachable sets for plan tracking and, at runtime, filters the reinforcement learning policy’s continuous actions to the nearest safe plan without throttling learning speed. Compared to Model Predictive Control (MPC), CBF, and Safety-set methods, which provide strong online guarantees but require heavier per-step optimization, RTS offers computational efficiency through its offline precomputation strategy. Similarly, while Radial Basis Function (RBF) and Gaussian Process (GP)-based safe learning methods offer flexible uncertainty modeling, they often rely on probabilistic guarantees and incur expensive computational costs. this approach relies on a reasonably accurate system model and can be less adaptive when dynamics or obstacles change significantly during operation.
@@
Model predictive shielding (MPS) [banerjee2024dynamic] provide a model-predictive shield around any RL policy: they forecast violations and replace unsafe actions with safe, goal-directed recoveries, enabling zero-violation training and deployment. CBF-Guided MPC [liu2025flexible] embeds barrier constraints in MPC and tunes them online for proactive, explainable hard-safety during tracking/avoidance at the cost of online optimization and careful model or CBF design. These MPC-based methods trade offline precomputation for online optimization to improve task progress with interpretable constraints.
@@
Implicit Safe Set Algorithm [zhao2024implicit]. A model-free safe control method that queries a black-box dynamics (e.g., digital twin) to synthesize a safety index (barrier certificate) and a corresponding safe control law, guaranteeing forward invariance and finite-time convergence. Data-Driven Permissible Safe Control with Barrier Certificates [mazouz2024data] learns unknown stochastic dynamics via Gaussian Processes and constructs piecewise stochastic barrier functions to compute a maximal set of permissible strategies that satisfy a target probabilistic safety level; as data increases, the permissible set grows toward the known-model benchmark while maintaining safety.
@@
Advantage-Based Intervention [wagener2021safe] achieves safety by online interventions and surrogate MDP training (requires a safe backup policy/advantage estimates), whereas AlwaysSafe [simao2021alwayssafe] achieves safety by a model-based abstraction of safety dynamics (requires a correct abstraction); both enable safe data collection for learning.
@@
Here we group learning-based methods by what they learn and how they ensure safety. Lyapunov actor–critic [16] learns policy weights under an explicit Lyapunov constraint, and executes through an adaptive impedance layer for compliant, stability-guaranteed motion. Safety-Aware Unsupervised Skill Discovery [22] learns a skill policy and a safety-critic that estimates failure probability, yielding intrinsically safe skills for downstream tasks. Variable-Impedance Skill Learning [23] learns a latent skill prior/decoder, a latent policy, and impedance as part of the action, achieving safety via compliant interaction and bounded contact forces. Impedance-learning adaptive control for HRI [9531394] learns impedance gains, feedforward interaction force, dynamics parameters, and robust Sliding Mode Control (SMC) gain, with Lyapunov proofs of bounded tracking. Impedance iterative-learning SMC [5] updates impedance and SMC gains across trials, combining compliance and robustness guarantees for non-repetitive tasks.
@@
3.1.2 Safe Execution (after contact).
@@
Safe execution is crucial in contact-rich robotic tasks, as robots must interact not only with complex and uncertain environments but often also in close proximity to humans [ajoudani2018progress]. During these interactions, any unsafe behavior can result in damage to the robot, the manipulated objects, or even pose risks to human safety. Ensuring safe execution means that the robot’s actions consistently respect physical constraints [141], avoid hazardous situations, and maintain robust performance despite uncertainties or disturbances. This is especially important in real-world deployments, where unpredictable events and unmodeled dynamics are common. Therefore, developing methods that guarantee safety during execution is essential for the reliable and trustworthy operation of robots in contact-rich scenarios.
@@
A widely adopted and effective approach to enhance safe execution in contact-rich robotic tasks is the use of compliant controllers [317, hogan1985impedance, 24, 23, samadikhoshkho_review_2025]. These controllers enable robots to interact safely with their environment and with humans by allowing controlled flexibility and adaptability in response to external forces. Among the most common compliant control strategies are impedance control and admittance control. Impedance control regulates the dynamic relationship between the robot’s motion (position or velocity) and the forces it experiences, effectively making the robot behave like a virtual mass-spring-damper system [hogan1985impedance]. This allows the robot to absorb impacts and maintain stable contact with uncertain or moving surfaces. Admittance control, on the other hand, measures external forces and commands the robot to move accordingly, making it suitable for tasks where the environment imposes significant forces on the robot [232, 205]. Both approaches are fundamental in ensuring that robots can safely and robustly perform manipulation tasks involving physical contact, reducing the risk of damage or injury during interaction.
@@
One key challenge in ensuring safe execution is properly tuning controller parameters (such as stiffness, damping, and inertia) to suit different task scenarios and contact conditions [calanca2015review]. Incorrect parameter settings can lead to unsafe behaviors, such as excessive contact forces, instability, or failure to accomplish the task. Traditionally, these parameters are manually tuned based on expert knowledge and trial and error, which is time-consuming and may not generalize well across diverse tasks or environments. Learning-based methods offer a promising solution by automatically adapting controller parameters through data-driven optimization, enabling robots to achieve both safety and high performance in varying and uncertain contact-rich scenarios [cruz2024learning, 10705419].

### v2

@@ 2.2.2 Stability and Invariance: Lyapunov and CBF Views
A second axis treats safety as stability and set invariance. Lyapunov methods certify convergence or boundedness and can imply safety when safe sets are invariant sublevel sets of a Lyapunov function [133, 145]. Control Barrier Functions (CBFs) enforce forward invariance of safety sets online, minimally modifying actions to prevent constraint violations; they have seen increasing use in contact-aware robotics and safe RL [8, 93, 177, 171]. End-to-end safe RL can embed barrier-based constraints into training or runtime filtering [48, 33]. Here, safety means remaining inside certified invariant sets while preserving task feasibility.
@@ 2.2.3 Constraint Satisfaction and Reachability
Safety can be defined as the satisfaction of state, input, and contact constraints (e.g., joint/velocity/torque limits, friction cones, force envelopes). Safe MPC enforces constraints over a horizon with robust tubes or uncertainty sets [101, 217]. Hamilton–Jacobi reachability and safety filters compute (or approximate) safe sets and veto actions that would exit them [17, 314, 259]. In the constraint-centric view, safety is the persistent satisfaction of physically meaningful bounds throughout approach, impact, sliding, and re-grasp phases.
@@ 2.2.4 Risk- and Data-Driven Safety
Data-driven definitions quantify safety via risk-sensitive objectives, recovery zones, or auxiliary critics that predict and penalize unsafe outcomes. Recovery RL separates exploration from recovery to reduce unsafe excursions [301]; safety critics or pessimistic objectives downweight risky actions under uncertainty [284, 92, 39]. Recent work provides theoretical formulations (e.g., safe value functions, shielding and generalization guarantees) to connect statistical learning with formal safety [207, 110]. In this axis, safety is a probabilistic guarantee or risk bound over contact outcomes under imperfect models and noisy sensing.
@@ 2.2.5 Human-Centered and Standard-Driven Safety
In human–robot collaboration and assistance, safety must address physical limits (contact pressure, speed-and-separation, ergonomics) and perceived safety (predictability, comfort, trust) [6, 73]. Standards such as ISO 10218 and ISO/TS 15066 codify limits and processes, shaping acceptable forces, speeds, and safety-rated monitoring [44, 149]. Perception- and policy-level safeguards further bound human exposure to risk in collaborative and assistive tasks [387, 68, 369, 290]. Along this axis, safety extends beyond robot/object protection to human well-being and acceptability.
2.2.6 Semantic Safety with Vision-Language(-Action) Models
@@
Semantic safety concerns aligning task intent, environment context, and contact constraints with what the robot actually executes, so that language-/vision-conditioned policies do not trigger unsafe contact behaviors. In contact-rich settings, this means grounding task semantics (objects, affordances, no-touch zones, insertion axes, force limits) and safety rules (limits, PPE, human proximity) into actionable constraints and monitors for execution [362].
@@
Modern VLM/VLA systems can (i) parse natural-language goals and safety rules, (ii) ground them in the sensed scene (objects, surfaces, grasp points) [342, 388, 367, 303], and (iii) parameterize low-level compliant controllers or skill libraries with safety-aware references (pose/force/impedance) [307, 393, 328] rather than direct torques. Typical architectures use a hierarchical split: a language–vision planner selects certified skills/options; mid-level skills expose safety-relevant parameters; low-level safety enforcers (impedance, CBF/MPC, rate limiters) guarantee contact stability and constraint satisfaction during execution [328, 58, 191, 370].
2.3 Problem Formulation and Learning Paradigm
2.3.1 Problem Setting
@@
Contact-rich manipulation arises across assembly, insertion, forceful surface interaction, and human–robot collaboration. We consider a general formulation that does not commit to a single mathematical model. Tasks are described by (i) a state representation $x_{t}$ that includes robot kinematics/dynamics, latent contact modes, and measured or estimated wrenches; (ii) a control signal $u_{t}$ that can be low-level torques/velocities or higher-level references such as force targets and impedance/admittance gains; and (iii) safety and feasibility specifications involving force/torque envelopes, friction and approach cones, motion limits, and human-centric constraints [6, 41, 362]. Depending on assumptions and requirements, the underlying decision model may be an unconstrained Markov decision process (MDP) or a partially observable Markov decision process (POMDP), a constrained variant (e.g., constrained Markov decision process (CMDP)), or a hybrid-systems/control-theoretic problem with explicit contact transitions. In all cases, safety is operationalized through limits on contact forces/pressures and motion, potentially augmented by risk- or semantics-derived rules [362, 367, 14, 307].
@@
Observability and abstraction are key design choices. When direct force/torque sensing is available, controllers can track force references and tighten safety envelopes; otherwise, observers or learned estimators provide wrench/contact-state cues for policy inputs and online monitoring [41]. Action abstraction often improves safety and sim-to-real transfer: outputting pose/force targets or impedance parameters instead of raw torques yields compliant interaction and clearer hooks for runtime safety mechanisms [41, 6, 362]. Finally, semantics (VLM/VLA) supply high-level task structure and safety rules that must be grounded into formal constraints and checked during execution [367, 14, 307, 82, 91].
Fig. 3 provides an overview of this survey. We begin by presenting the background and key definitions in Sec.3 then organize existing approaches from multiple perspectives, covering task characteristics, sensing and policy modalities, data acquisition strategies, simulation environments and benchmarks (Sec.4) and promising future directions (Sec. 5).
[…]
Figure 3: Structure of this survey. We first introduce the background and definitions in Sec. 2, then categorize existing methods from different perspectives in Sec. 3, including task characteristics, sensing and policy modalities, data acquisition strategies, simulation environments and benchmarks (Sec. 3.5), safety evaluation metrics, safety abstraction levels, and safety enforcement spaces. Finally, we discuss open challenges (Sec. 4) and future directions (Sec. 5).
@@ 2.3.2 Learning Paradigm
We summarize five main learning paradigms and how they interface with safety in contact-rich manipulation.
@@
Reinforcement learning: RL learns policies for variable-impedance forceful skills and insertions, often using action spaces that output force targets or stiffness, damping to preserve compliant interaction. Safety is enhanced by conservative objectives, recovery strategies, and shielded execution (e.g., barrier, reachability, MPC layers) [374, 248, 285, 373, 173, 159]
@@
Imitation learning: In IL, initial policies of contact sequencing and force behavior (approach, touch, slide) are learned from demonstrations. It is common to refine the policy by RL or supervised updates; hybrid trajectory-force learning is prevalent in insertion and assembly settings. IL shares the same safety-friendly action abstractions and leverages compliant controllers during execution [193, 325, 198]. In this survey, we focus on methods with a notable online learning component, and direct the readers to recent surveys for pure IL methods [310].
@@
Learning for model-based and control-theoretic methods: Impedance and admittance regulation, robust and adaptive force controllers, MPC with contact constraints, and barrier, reachability filters operationalize limits on contact force, torque, friction cones, and etc. Learning augments these by identifying dynamics, adapting gains, or shaping feasible sets, yielding transparent enforcement at runtime [10, 171, 392, 238, 10, 346]
@@
Foundation model-based methods: Foundation models for robotics leverage large-scale vision-language (and action) pretraining to parse open-ended goals, infer affordances, and impose safety-aware constraints before execution. In contact-rich settings, these methods are typically embedded in a plan-parameterize-enforce stack: (i) a VLM/VLA module interprets natural-language goals and scene context to propose task sketches and safety rules (e.g., force threshold, no-touch regions, speed limitation); (ii) the semantic plan is compiled into formal constraints and parameterized skill references (pose, force, impedance) rather than raw torques to preserve compliant interaction [342]; and (iii) low-level safety layers (e.g., MPC/CBF filters, rate limiters) execute with verifiable bounds on contact forces and approach geometry [367, 14, 307]. Data and evaluation rely on open toolchains and safety-oriented benchmarks to quantify violation rates, refusal/recovery behavior, and generalization under distribution shift [308, 123, 82, 91]. Practically, safety guardrails for foundation-model pipelines include structured prompting and tool schemas, text-to-constraint grounding with pre-execution checks, conservative default limits (force/velocity caps), and “plan-then-filter” action projection through certified controllers; these mitigations address hallucination and mis-grounding that could otherwise lead to overloads at contact [367, 307, 388].
@@
Hybrid integration Planning and skill graphs (potentially guided by VLM/VLA) provide structure; learned policies output pose/force/impedance parameters; safety layers project actions to satisfy certified constraints. Semantic modules further align high-level intent and safety rules with the certified execution stack [82, 91, 307, 388]. Evaluation commonly reports violations, near misses, and generalization using standardized suites and simulators [308, 243, 123].
3 Categorization of Methods from Different Perspectives
@@ 3.1 Safe Learning for Contact-Rich Tasks
In this section, we categorize existing safe learning approaches for contact-rich robot tasks based on the phase: safe exploration (learning before contact) and safe execution (learning during contact) and provable safety methods, and discuss representative methods in each category. This categorization helps clarify the distinct challenges and strategies associated with each phase, as well as their implications for safety in contact-rich manipulation.
3.1.1 Safe Exploration (safe learning before contact).
@@
Safe human-robot cooperation (HRC) is a major application branch within the broader class of contact-rich tasks, and it has traditionally been approached through path planning algorithms such as rapidly-exploring random tree star (RRT[50] for dynamic human-robot systems, or artificial potential field (APF) [246] and virtual fixtures [34] for real-time adjustment of the robot’s trajectory. However, these methods often yield slow trajectory computation and limited generalization, since they require task-specific tuning, assume simplified environments, and adapt poorly to dynamic or unstructured settings [34].
Recent approaches enforce safety constraints via control barrier functions (CBFs), where quadratic programs compute minimal trajectory deviations to maintain safety [109, 76, 59]. Extensions to uncertain CBFs (U-CBFs) handle environmental uncertainty [63, 161], though worst-case assumptions often yield conservative designs. Stochastic frameworks using chance constraints offer less conservative alternatives [318, 306], but remain computationally intractable for most HRC applications due to non-convex feasible regions [5]. Complementary work examines risk factors in HRC [26, 118, 187, 201]. While these approaches can better inform the tolerable risk-bound for different applications, a risk tunable CBF can solve the limitation [264].
@@
Across these reachability-based methods, a common theme is to shift most of the computational burden into the offline construction of trajectory-parameterized reachable sets based on a reasonably accurate model. Reachability-based Trajectory Safeguard (RTS) [260] is an offline–online safety layer for continuous-control RL that precomputes trajectory-parameterized forward reachable sets and, at runtime, projects the RL agent’s chosen plan parameters to a nearest safe plan while preserving learning efficiency. A reachability-based manipulator planner  [105] that builds joint-space, trajectory-parameterized reachable sets offline and then uses them to constrain receding-horizon optimization.  [141] introduce a receding-horizon planner for mobile robots that performs offline reachability analysis over a continuum of trajectory parameters and then selects safe parameters online. Those approach rely on a reasonably accurate system model and can be less adaptive when dynamics or obstacles change significantly during operation.
@@
Model predictive shielding (MPS) [16] provide a model-predictive shield around any RL policy: they forecast violations and replace unsafe actions with safe, goal-directed recoveries, enabling zero-violation training and deployment. CBF-Guided MPC [178] embeds barrier constraints in MPC and tunes them online for proactive, explainable hard-safety during tracking/avoidance at the cost of online optimization and careful model or CBF design. These MPC-based methods trade offline precomputation for online optimization to improve task progress with interpretable constraints. Planner-augmented approaches such as MoPA-RL [347] can be viewed as a broader shield that uses classical motion planning (and reactive shaping) to intervene and generate collision-free long-horizon actions around a learned policy, improving exploration and execution safety without CBF-style hard guarantees.
RTS-style reachability safeguards shift burden offline, enabling fast runtime projection with hard safety so long as the precomputed reachable sets remain valid; their main limitation is reduced adaptivity under model mismatch or rapidly changing contact dynamics. In contrast, MPS and CBF-guided MPC place more computation online: they can respond to changing constraints and goals within a receding horizon, but require repeated optimization and careful dynamics/constraint modeling, with feasibility and real-time performance becoming the bottleneck in high-dimensional HRC. Practically, reachability filters suit fixed platforms with stable dynamics and tight real-time budgets, whereas MPC/CBF variants suit environments where constraints and intent change quickly and online compute is available.
@@
Implicit Safe Set Algorithm [384]. A model-free safe control method that queries a black-box dynamics (e.g., digital twin) to synthesize a safety index (barrier certificate) and a corresponding safe control law, guaranteeing forward invariance and finite-time convergence. Data-Driven Permissible Safe Control with Barrier Certificates [211] learns unknown stochastic dynamics via Gaussian Processes and constructs piecewise stochastic barrier functions to compute a maximal set of permissible strategies that satisfy a target probabilistic safety level; as data increases, the permissible set grows toward the known-model benchmark while maintaining safety. Together with safe exploration methods [279, 349, 156], these approaches illustrate how pre-deployment safe learning can constrain exploration to certified-safe behaviors before execution.
@@
Advantage-Based Intervention [315] achieves safety by online interventions and surrogate MDP training (requires a safe backup policy/advantage estimates), whereas AlwaysSafe [270] achieves safety by a model-based abstraction of safety dynamics (requires a correct abstraction); both enable safe data collection for learning.
@@
Here we group learning-based methods by what they learn and how they ensure safety. Lyapunov actor–critic [165] learns policy weights under an explicit Lyapunov constraint, and executes through an adaptive impedance layer for compliant, stability-guaranteed motion. Safety-Aware Unsupervised Skill Discovery [138] learns a skill policy and a safety-critic that estimates failure probability, yielding intrinsically safe skills for downstream tasks. Variable-Impedance Skill Learning [350] learns a latent skill prior/decoder, a latent policy, and impedance as part of the action, achieving safety via compliant interaction and bounded contact forces. Impedance-learning adaptive control for HRI [262] learns impedance gains, feedforward interaction force, dynamics parameters, and robust Sliding Mode Control (SMC) gain, with Lyapunov proofs of bounded tracking. Impedance iterative-learning SMC [345] updates impedance and SMC gains across trials, combining compliance and robustness guarantees for non-repetitive tasks. Model-accelerated maximum-entropy RL for high-precision assembly [385] learns a stochastic policy executed through an impedance controller and leverages a GP dynamics model for sample-efficient training, achieving practical safety via compliant interaction.
@@
Since the literature lacks a unified benchmark that evaluates different safety paradigms on the same contact-rich task under matched conditions, the comparisons in this section are necessarily synthesis-based: we distill the reported assumptions, guarantees, and practical constraints from the above studies into a common set of criteria. To provide a unified lens for analyzing trade-offs across method families, we compare them using six recurring axes: (A1) guarantee strength/type (hard invariance/reachability vs. probabilistic vs. empirical), (A2) model reliance (known dynamics vs. approximate/black-box vs. data-driven), (A3) compute placement (offline precomputation vs. online optimization/filtering), (A4) conservatism and risk tuning (worst-case bounds vs. chance levels vs. tunable safety margins), (A5) scalability and real-time viability (state dimension, horizon length, contacts/multi-agent), and (A6) data/sample efficiency (how safety constraints affect exploration and learning speed). In figures 4, we summarize these axes using hexagonal ratings for representative methods.
Figure 4: Hexagonal ratings across six comparison axes (higher values indicate stronger relative performance on a 1–5 scale). (A1) Guarantee strength: 1 = empirical/heuristic only, 5 = hard safety (invariance/reachability-style). (A2) Model independence: 1 = requires an accurate known model, 5 = model-free / black-box / learned abstraction. (A3) Online efficiency: 1 = heavy online optimization, 5 = lightweight runtime (most computation offline). (A4) Conservatism / tunability: 1 = highly conservative with few tuning knobs, 5 = explicitly risk-tunable and typically less conservative in practice. (A5) Scalability / real-time viability: 1 = struggles in high-dimensional contact-rich or multi-agent settings, 5 = scales well and runs in real time. (A6) Data efficiency (safe learning): 1 = data-hungry, 5 = requires little data to ensure safety or supports efficient safe data collection.
Across the six axes (A1–A6), the literature reveals a recurring trade-off between guarantee strength and adaptability/compute. Classical planning and reactive shaping are intuitive but often provide weaker or indirect guarantees and require task-specific tuning, limiting scalability in dynamic HRC. CBF-style methods offer explicit, interpretable safety enforcement with strong guarantees when modeling assumptions hold, but can become conservative under uncertainty (e.g., worst-case U-CBFs) or computationally heavy when using less conservative probabilistic constraints. Learning-based compliant skills improve adaptability to complex contact phenomena and can achieve practical safety via compliance and stability arguments, but their effective safety depends on training coverage and may trade formal hardness for data requirements. Reachability-based safeguards deliver the strongest hard-safety behavior with predictable runtime by shifting computation offline, yet rely on accurate models and can be less adaptive under mismatch or rapidly changing environments. MPC/shielding approaches sit between these extremes, improving adaptivity through online prediction and constraint handling but incurring repeated optimization costs and feasibility concerns in high-dimensional contact-rich settings. Safe-set/certificate construction and intervention/abstraction-based safe exploration further enable pre-deployment or training-time safety, but their conservatism and scalability are governed by certificate fidelity, backup/abstraction correctness, and how often safety mechanisms override exploration.
3.1.2 Safe Execution (post contact stage).
@@
Safe execution is crucial in contact-rich robotic tasks, as robots must interact not only with complex and uncertain environments but often also in close proximity to humans [6]. During these interactions, any unsafe behavior can result in damage to the robot, the manipulated objects, or even pose risks to human safety. Ensuring safe execution means that the robot’s actions consistently respect physical constraints [96], avoid hazardous situations, and maintain robust performance despite uncertainties or disturbances. This is especially important in real-world deployments, where unpredictable events and unmodeled dynamics are common. Therefore, developing methods that guarantee safety during execution is essential for the reliable and trustworthy operation of robots in contact-rich scenarios.
@@
A widely adopted and effective approach to enhance safe execution in contact-rich robotic tasks is the use of compliant controllers [222, 103, 261, 350, 253]. These controllers enable robots to interact safely with their environment and with humans by allowing controlled flexibility and adaptability in response to external forces. Among the most common compliant control strategies are impedance control and admittance control. Impedance control regulates the dynamic relationship between the robot’s motion (position or velocity) and the forces it experiences, effectively making the robot behave like a virtual mass-spring-damper system [103]. This allows the robot to absorb impacts and maintain stable contact with uncertain or moving surfaces. Admittance control, on the other hand, measures external forces and commands the robot to move accordingly, making it suitable for tasks where the environment imposes significant forces on the robot [289, 255]. Both approaches are fundamental in ensuring that robots can safely and robustly perform manipulation tasks involving physical contact, reducing the risk of damage or injury during interaction.
@@
One key challenge in ensuring safe execution is properly tuning controller parameters (such as stiffness, damping, and inertia) to suit different task scenarios and contact conditions [41]. Incorrect parameter settings can lead to unsafe behaviors, such as excessive contact forces, instability, or failure to accomplish the task. Traditionally, these parameters are manually tuned based on expert knowledge and trial and error, which is time-consuming and may not generalize well across diverse tasks or environments. Learning-based methods offer a promising solution by automatically adapting controller parameters through data-driven optimization, enabling robots to achieve both safety and high performance in varying and uncertain contact-rich scenarios [57, 388].